OpenStudy (anonymous):

ABC is a triangle whose three sides are different. Its perimeter is 20 cm. When the longest side of the triangle is shortened by 2 cm, an isosceles triangle is formed. If that side is further shortened by 3 cm, another isosceles triangle is formed. Find the lengths of the sides of the original triangle ABC. @Peter14 @satellite73

OpenStudy (anonymous):

try to express the conditions as a system of equations. what do you have so far?

OpenStudy (anonymous):

let x cm be the length of the longest side. (x-2)+(x-3)+x=20 i think it's wrong..

OpenStudy (anonymous):

but my answer is wrong..

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

o.O

OpenStudy (anonymous):

"if that side is further shortened by 3 cm"

OpenStudy (anonymous):

sorry for missing that the first time

OpenStudy (anonymous):

ohh!!!

OpenStudy (anonymous):

(x-2)+(x-2-3)+x=20

OpenStudy (anonymous):

x=9

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

then the second longest side is 9-2=7cm

OpenStudy (anonymous):

the shortest side is 9-2-3=4cm

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thanks